# thanks: Present address: Google Quantum AI, Santa Barbara, CAthanks: Present address: Google Quantum AI, Santa Barbara, CA

### v1

@@ Abstract
Hilbert space dimension is a key resource for quantum information processing. A large Hilbert space is not only an essential requirement for quantum error correction, but it can also be advantageous for realizing gates and algorithms more efficiently. There has thus been considerable experimental effort in recent years to develop quantum computing platforms using qudits (d-dimensional quantum systems with d >2) as the fundamental unit of quantum information. Just as with qubits, quantum error correction of these qudits will be necessary in the long run, but to date error correction of logical qudits has not been demonstrated experimentally. Here we report the experimental realization of error-corrected logical qutrits (d=3) and ququarts (d=4) by employing the Gottesman-Kitaev-Preskill (GKP) bosonic code in a circuit QED architecture. Using a reinforcement learning agent, we optimize the GKP qutrit (ququart) as a ternary (quaternary) quantum memory and achieve beyond break-even error correction with a gain of 1.82 ± 0.03 (1.87 ± 0.03). This work represents a new way of leveraging the large Hilbert space of a harmonic oscillator for hardware-efficient quantum error correction.
@@
On the other hand, embracing these qudits can enable more efficient distillation of magic states [campbell_enhanced_2014], synthesis of gates [ralph_efficient_2007, fedorov_implementation_2012], compilation of algorithms [bocharov_factoring_2017, kiktenko_scalable_2020, gokhale_asymptotic_2019], and simulation of high-dimensional quantum systems [meth_simulating_2023, sawaya_resource-efficient_2020]. For these reasons, there has been considerable experimental effort in recent years to develop qudit-based platforms for quantum computing, for example using donor spins in silicon [asaad_coherent_2020, fernandez_de_fuentes_navigating_2024, yu_creation_2024], ultracold atoms and molecules [vilas_optical_2024, lindon_complete_2023, smith_quantum_2013, chaudhury_quantum_2007], optical photons [chi_programmable_2022, giovannini_characterization_2013, bent_experimental_2015, yoshikawa_heralded_2018, reimer_high-dimensional_2019, babazadeh_high-dimensional_2017, wang_multidimensional_2018, kues_-chip_2017, zhang_spinorbit_2022], superconducting circuits [champion_multi-frequency_2024, goss_high-fidelity_2022, morvan_qutrit_2021, neeley_emulation_2009, nguyen_empowering_2024, roy_two-qutrit_2023, roy_synthetic_2024, svetitsky_hidden_2014, wang_systematic_2024], trapped ions [hrmo_native_2023, leupold_sustained_2018, low_control_2023, ringbauer_universal_2022, senko_realization_2015], and vacancy centers [adambukulam_hyperfine_2024, singh_multi-photon_2022, soltamov_excitation_2019]. If qudits are to be useful in the long run, however, quantum error correction (QEC) will be necessary.
@@
Non-binary Quantum Memories
To evaluate the performance of our logical qudits as non-binary quantum memories, and establish the concept of QEC gain for qudits, we follow previous work [sivak_real-time_2023] and use the average channel fidelity $\mathcal{F}_{d}(\mathcal{E},I)$, which quantifies how well a channel $\mathcal{E}$ realizes the identity $I$ [nielsen_simple_2002]. Although $\mathcal{F}_{d}$ will have a non-exponential time evolution in general, it can always be expanded to short times $dt$ as $\mathcal{F}_{d}(\mathcal{E},I)\approx 1-\frac{d-1}{d}\Gamma dt$, where $\Gamma$ is the effective decay rate of the channel $\mathcal{E}$ at short times [noauthor_see_nodate]. This rate $\Gamma$ enables us to compare different decay channels on the same footing. In particular, we want to compare the decay rate $\Gamma^{\mathrm{logical}}_{d}$ of our logical qudit to $\Gamma^{\mathrm{physical}}_{d}$ of the best physical qudit in our system. We define the QEC gain as their ratio $G_{d}=\Gamma^{\mathrm{physical}}_{d}/\Gamma^{\mathrm{logical}}_{d}$, and the break-even point is when this gain is unity.
@@
Figure 2: Logical GKP Qutrit. a, State preparation of qutrit Pauli eigenstates $|P_{0}\rangle_{3}$ with $\Delta=0.32$. b, Circuit for measuring a qutrit in the basis of Pauli operator $P_{3}$ using an ancilla qubit, where $\theta_{0}=2\arctan(1/\sqrt{2})$. The first measurement distinguishes between the state $|P_{0}\rangle_{3}$ and the subspace $\{|P_{1}\rangle_{3},|P_{2}\rangle_{3}\}$, while the second distinguishes between $|P_{1}\rangle_{3}$ and $\{|P_{0}\rangle_{3},|P_{2}\rangle_{3}\}$. The Bloch spheres depict the trajectories taken by the ancilla when the qutrit is in each Pauli eigenstate. c, Backaction of the GKP qutrit Pauli measurement in the $Z_{3}$ basis, applied to the maximally mixed qutrit state. d, Decay of qutrit Pauli eigenstates $|P_{0}\rangle_{3}$ under the optimized QEC protocol. The dashed black lines indicate a probability of $1/3$. The solid grey lines are exponential fits. From left to right we find $\gamma_{X_{0}}^{-1}=1153\pm 13\mathrm{\mu s}$, $\gamma_{Z_{0}}^{-1}=1120\pm 15\mathrm{\mu s}$, $\gamma_{XZ_{0}}^{-1}=743\pm 10\mathrm{\mu s}$, and $\gamma_{X^{2}Z_{0}}^{-1}=727\pm 11\mathrm{\mu s}$.
@@
Figure 3: Logical GKP Ququart. a, State preparation of ququart Pauli eigenstates $|P_{0}\rangle_{4}$ and parity state $|+,0\rangle_{4}$ with $\Delta=0.32$. b, Circuit for measuring a ququart in the basis of Pauli operator $P_{4}$ using an ancilla qubit. The first measurement distinguishes between the even and odd states $|P_{\mathrm{even/odd}}\rangle_{4}$, and the second measurement distinguishes between the remaining two states. The Bloch spheres depict the trajectories taken by the ancilla when the ququart is in each Pauli eigenstate $|P_{n}\rangle_{4}$. c, Backaction of the GKP ququart Pauli measurement in the $Z_{4}$ basis, applied to the maximally mixed ququart state. d, Circuit for measuring a ququart in the parity basis {|±$\{|\pm,m\rangle_{4}:m=0,1\}$, where $X_{4}^{2}|\pm,m\rangle_{4}=\pm|\pm,m\rangle_{4}$ and $Z_{4}^{2}|\pm,m\rangle_{4}=(-1)^{m}|\pm,m\rangle_{4}$. The first measurement determines the eigenvalue of $X_{4}^{2}$, and the second measurement determines the eigenvalue of $Z_{4}^{2}$. e, Backaction of the GKP ququart parity measurement, applied to the maximally mixed ququart state. f, Decay of ququart Pauli eigenstates $|P_{0}\rangle_{4}$ and parity state $|+,0\rangle_{4}$ under the optimized quantum error correction protocol. The dashed black lines indicate a probability of $1/4$. The solid grey lines are exponential fits. From left to right we find $\gamma_{X_{0}}^{-1}=840\pm 8\mathrm{\mu s}$, $\gamma_{Z_{0}}^{-1}=836\pm 9\mathrm{\mu s}$, $\gamma_{XZ_{0}}^{-1}=519\pm 6\mathrm{\mu s}$, $\gamma_{X^{2}Z_{0}}^{-1}=507\pm 9\mathrm{\mu s}$, $\gamma_{X^{3}Z_{0}}^{-1}=571\pm 7\mathrm{\mu s}$, $\gamma_{XZ^{2}_{0}}^{-1}=562\pm 9\mathrm{\mu s}$, and $\gamma_{+,0}^{-1}=607\pm 8\mathrm{\mu s}$.
@@
We measure the GKP ququart in Pauli basis $P_{4}$ using the circuit shown in Fig. 3b. As with the qutrit, the generalized ancilla-qubit-controlled Pauli operators $CP_{4}=|g\rangle\langle g|I_{4}+|e\rangle\langle e|P_{4}$ are realized via ECD gates (see Methods). Note that the measurement circuit is simpler than the case of the qutrit because it is easier to sequentially split the $d=4$ Hilbert space of a ququart in half symmetrically via binary measurements of an ancilla qubit. The first measurement distinguishes between the even subspace $\{|P_{0}\rangle,|P_{2}\rangle\}$ and odd subspace $\{|P_{1}\rangle,|P_{3}\rangle\}$ by measuring whether $P_{4}^{2}=\pm 1$, while the second measurement distinguishes between the remaining two states by measuring $P_{4}=\pm 1$ (if in the even subspace) or $P_{4}=\pm i$ (if in the odd subspace). To verify that this circuit realizes the desired projective measurement, we prepare the maximally mixed state of the GKP ququart $\rho_{4}^{\mathrm{mix}}$, measure in the $Z_{4}$ basis, and perform Wigner tomography of the cavity postselected on the four measurement outcomes. The results of this measurement are shown in Fig. 3c (see [noauthor_see_nodate] for measurements in the other Pauli bases).
@@ Discussion
With a smaller $\Delta$, our logical GKP qudits have more energy, which should amplify the rate of cavity photon loss, and have information stored further out in phase space, which should amplify the effects of cavity dephasing. To corroborate this, we simulate our optimal QEC protocols and isolate the relative contributions of different physical errors to our overall logical error rates [noauthor_see_nodate]. We find that the three largest sources of logical errors are transmon bit-flips, whose relative contribution decreases as $d$ increases, cavity photon loss, whose relative contribution increases as $d$ increases, and cavity dephasing, which is the dominant source of error and whose relative contribution increases as $d$ increases. Since our cavity dephasing is primarily due to the thermal population $n_{\mathrm{th}}=2.2\pm 0.1\%$ of the transmon [reagor_quantum_2016], the lifetimes of our logical GKP qudits could be substantially improved by either reducing $n_{\mathrm{th}}$ or using an ancilla that can be actively decoupled from the cavity when not in use [rosenblum_fault-tolerant_2018, koottandavida_erasure_2024, ding_quantum_2024].
In summary, we have demonstrated QEC of logical qudits with $d>2$, which represents a milestone achievement in the development of qudits for useful quantum technologies. Moreover, we have beaten the break-even point for QEC of quantum memories, a result few other experiments have accomplished [ofek_extending_2016, ni_beating_2023, sivak_real-time_2023, acharya_quantum_2024]. This work builds on the promise of hardware efficiency offered by bosonic codes [cai_bosonic_2021, gertler_protecting_2021, ofek_extending_2016, ni_beating_2023], and represents a new way of leveraging the large Hilbert space of an oscillator. In exchange for a modest reduction in lifetime, we gain access to more logical quantum states in a single physical system. This could enable more efficient compilation of gates [ralph_efficient_2007, fedorov_implementation_2012] and algorithms [bocharov_factoring_2017, kiktenko_scalable_2020, gokhale_asymptotic_2019], novel techniques for quantum communication [schmidt_error-corrected_2024] and transduction [wang_passive_2024], and advantageous strategies for concatenation into an external multi-qudit code [gottesman_stabilizer_1997, gottesman_fault-tolerant_1998, campbell_magic-state_2012, campbell_enhanced_2014]. With the realization of bosonic logical qudits we also establish a platform for concatenating codes internally: by embedding a logical qubit within a bosonic logical qudit [gottesman_encoding_2001, pirandola_minimal_2008, cafaro_quantum_2012, omanakuttan_fault-tolerant_2024, gross_designing_2021, gross_hardware-efficient_2024], multiple layers of error correction could be implemented inside a single oscillator.
@@ Methods
Ethics declarations: LF is a founder and shareholder of Quantum Circuits Inc. (QCI). The remaining authors declare no competing interests.

### v2

@@ Abstract
Hilbert space dimension is a key resource for quantum information processing. A large Hilbert space is not only an essential requirement for quantum error correction, but it can also be advantageous for realizing gates and algorithms more efficiently. There has thus been considerable experimental effort in recent years to develop quantum computing platforms using qudits (d-dimensional quantum systems with d >2) as the fundamental unit of quantum information. Just as with qubits, quantum error correction of these qudits will be necessary in the long run, but to date error correction of logical qudits has not been demonstrated experimentally. Here we report the experimental realization of an error-corrected logical qutrit (d=3) and ququart (d=4) by employing the Gottesman-Kitaev-Preskill (GKP) bosonic code. Using a reinforcement learning agent, we optimize the GKP qutrit (ququart) as a ternary (quaternary) quantum memory and achieve beyond break-even error correction with a gain of 1.82 ± 0.03 (1.87 ± 0.03). This work represents a new way of leveraging the large Hilbert space of a harmonic oscillator for hardware-efficient quantum error correction.
@@
On the other hand, embracing these qudits can enable more efficient distillation of magic states [campbell_enhanced_2014], synthesis of gates [ralph_efficient_2007, fedorov_implementation_2012], compilation of algorithms [bocharov_factoring_2017, kiktenko_scalable_2020, gokhale_asymptotic_2019, Chu2023], and simulation of high-dimensional quantum systems [meth_simulating_2023, sawaya_resource-efficient_2020]. For these reasons, there has been considerable experimental effort in recent years to develop qudit-based platforms for quantum computing, for example using donor spins in silicon [asaad_coherent_2020, fernandez_de_fuentes_navigating_2024, yu_creation_2024], ultracold atoms and molecules [vilas_optical_2024, lindon_complete_2023, smith_quantum_2013, chaudhury_quantum_2007], optical photons [chi_programmable_2022, giovannini_characterization_2013, bent_experimental_2015, yoshikawa_heralded_2018, reimer_high-dimensional_2019, babazadeh_high-dimensional_2017, wang_multidimensional_2018, kues_-chip_2017, zhang_spinorbit_2022], superconducting circuits [champion_multi-frequency_2024, goss_high-fidelity_2022, morvan_qutrit_2021, neeley_emulation_2009, nguyen_empowering_2024, roy_two-qutrit_2023, roy_synthetic_2024, svetitsky_hidden_2014, wang_systematic_2024], trapped ions [hrmo_native_2023, leupold_sustained_2018, low_control_2023, ringbauer_universal_2022, senko_realization_2015], and vacancy centers [adambukulam_hyperfine_2024, singh_multi-photon_2022, soltamov_excitation_2019]. If qudits are to be useful in the long run, however, quantum error correction (QEC) will be necessary.
@@
Characterizing Quantum Memories
To characterize the performance of our logical qudits as quantum memories, and establish the concept of QEC gain for qudits, we follow previous work [sivak_real-time_2023] and use the average channel fidelity $\mathcal{F}_{d}(\mathcal{E},I)$, which quantifies how well a channel $\mathcal{E}$ realizes the identity $I$ [nielsen_simple_2002]. Although $\mathcal{F}_{d}$ will have a non-exponential time evolution in general, it can always be expanded to short times $dt$ as $\mathcal{F}_{d}(\mathcal{E},I)\approx 1-\frac{d-1}{d}\Gamma dt$, where $\Gamma$ is the effective decay rate of the channel $\mathcal{E}$ at short times [noauthor_see_nodate]. This rate $\Gamma$ enables us to compare different decay channels on the same footing. In particular, we want to compare the decay rate $\Gamma^{\mathrm{logical}}_{d}$ of our logical qudit to $\Gamma^{\mathrm{physical}}_{d}$ of the best physical qudit in our system. We define the QEC gain as their ratio $G_{d}=\Gamma^{\mathrm{physical}}_{d}/\Gamma^{\mathrm{logical}}_{d}$, and the break-even point is when this gain is unity.
@@
Figure 2: Realization of a logical GKP qutrit. a, State preparation of qutrit Pauli eigenstates $|P_{0}\rangle_{3}$ with $\Delta=0.32$. b, Circuit for measuring a qutrit in the basis of Pauli operator $P_{3}$ using an ancilla qubit, where $\theta_{0}=2\arctan(1/\sqrt{2})$. The first measurement distinguishes between the state $|P_{0}\rangle_{3}$ and the subspace $\{|P_{1}\rangle_{3},|P_{2}\rangle_{3}\}$, while the second distinguishes between $|P_{1}\rangle_{3}$ and $\{|P_{0}\rangle_{3},|P_{2}\rangle_{3}\}$. The Bloch spheres depict the trajectories taken by the ancilla when the qutrit is in each Pauli eigenstate. c, Backaction of the qutrit Pauli measurement in the $Z_{3}$ basis, applied to the maximally mixed qutrit state. d, Decay of qutrit Pauli eigenstates $|P_{0}\rangle_{3}$ under the optimized QEC protocol. The dashed black lines indicate a probability of $1/3$. The solid grey lines are exponential fits. From left to right we find $\gamma_{X_{0}}^{-1}=1153\pm 13\mathrm{\mu s}$, $\gamma_{Z_{0}}^{-1}=1120\pm 15\mathrm{\mu s}$, $\gamma_{XZ_{0}}^{-1}=743\pm 10\mathrm{\mu s}$, and $\gamma_{X^{2}Z_{0}}^{-1}=727\pm 11\mathrm{\mu s}$.
@@
Figure 3: Realization of a logical GKP ququart. a, State preparation of ququart Pauli eigenstates $|P_{0}\rangle_{4}$ and parity state $|+,0\rangle_{4}$ with $\Delta=0.32$. b, Circuit for measuring a ququart in the basis of Pauli operator $P_{4}$ using an ancilla qubit. The first measurement distinguishes between the even and odd states $|P_{\mathrm{even/odd}}\rangle_{4}$, and the second measurement distinguishes between the remaining two states. The Bloch spheres depict the trajectories taken by the ancilla when the ququart is in each Pauli eigenstate. c, Backaction of the GKP ququart Pauli measurement in the $Z_{4}$ basis, applied to the maximally mixed ququart state. d, Circuit for measuring a ququart in the parity basis {|±$\{|\pm,m\rangle_{4}:m=0,1\}$, where $X_{4}^{2}|\pm,m\rangle_{4}=\pm|\pm,m\rangle_{4}$ and $Z_{4}^{2}|\pm,m\rangle_{4}=(-1)^{m}|\pm,m\rangle_{4}$. The first measurement determines the eigenvalue of $X_{4}^{2}$, and the second determines that of $Z_{4}^{2}$. e, Backaction of the GKP ququart parity measurement, applied to the maximally mixed ququart state. f, Decay of ququart Pauli eigenstates $|P_{0}\rangle_{4}$ and parity state $|+,0\rangle_{4}$ under the optimized QEC protocol. The dashed black lines indicate a probability of $1/4$. The solid grey lines are exponential fits. From left to right we find $\gamma_{X_{0}}^{-1}=840\pm 8\mathrm{\mu s}$, $\gamma_{Z_{0}}^{-1}=836\pm 9\mathrm{\mu s}$, $\gamma_{XZ_{0}}^{-1}=519\pm 6\mathrm{\mu s}$, $\gamma_{X^{2}Z_{0}}^{-1}=507\pm 9\mathrm{\mu s}$, $\gamma_{X^{3}Z_{0}}^{-1}=571\pm 7\mathrm{\mu s}$, $\gamma_{XZ^{2}_{0}}^{-1}=562\pm 9\mathrm{\mu s}$, and $\gamma_{+,0}^{-1}=607\pm 8\mathrm{\mu s}$.
@@
We measure the GKP ququart in Pauli basis $P_{4}$ using the circuit shown in Fig. 3b. As with the qutrit, the generalized ancilla-qubit-controlled Pauli operators $CP_{4}=|g\rangle\langle g|I_{4}+|e\rangle\langle e|P_{4}$ are realized via ECD gates (see Methods). Note that the measurement circuit is simpler than the case of the qutrit because it is easier to sequentially split the $d=4$ Hilbert space of a ququart in half via binary measurements of an ancilla qubit. The first measurement distinguishes between the even subspace $\{|P_{0}\rangle,|P_{2}\rangle\}$ and odd subspace $\{|P_{1}\rangle,|P_{3}\rangle\}$ by measuring whether $P_{4}^{2}=\pm 1$, while the second measurement distinguishes between the remaining two states by measuring $P_{4}=\pm 1$ (if in the even subspace) or $P_{4}=\pm i$ (if in the odd subspace). To verify that this circuit realizes the desired projective measurement, we prepare the maximally mixed state of the GKP ququart $\rho_{4}^{\mathrm{mix}}$, measure in the $Z_{4}$ basis, and perform Wigner tomography of the cavity postselected on the four measurement outcomes. The results of this measurement are shown in Fig. 3c (see [noauthor_see_nodate] for measurements in the other Pauli bases).
@@ Discussion
With a smaller $\Delta$, our logical GKP qudits have more energy and are more highly squeezed, which should amplify the rates of cavity photon loss and dephasing. To corroborate this, we simulate our optimal QEC protocols and isolate the relative contributions of different physical errors to our overall logical error rates [noauthor_see_nodate]. We find that the three largest sources of logical errors are transmon bit-flips, whose relative contribution decreases as $d$ increases, cavity photon loss, whose relative contribution increases as $d$ increases, and cavity dephasing, which is the dominant source of error and whose relative contribution increases as $d$ increases. Since our cavity dephasing is primarily due to the thermal population $n_{\mathrm{th}}=2.2\pm 0.1\%$ of the transmon [reagor_quantum_2016], the lifetimes of our logical GKP qudits could be substantially improved by either reducing $n_{\mathrm{th}}$ or using an ancilla that can be actively decoupled from the cavity when not in use [rosenblum_fault-tolerant_2018, koottandavida_erasure_2024, ding_quantum_2024].
In summary, we have demonstrated QEC of logical qudits with $d>2$, which represents a milestone achievement in the development of qudits for useful quantum technologies. Moreover, we have beaten the break-even point for QEC of quantum memories, a result few other experiments have accomplished [ofek_extending_2016, ni_beating_2023, sivak_real-time_2023, acharya_quantum_2024]. This work builds on the promise of hardware efficiency offered by bosonic codes [cai_bosonic_2021, gertler_protecting_2021, ofek_extending_2016, ni_beating_2023], and represents a new way of leveraging the large Hilbert space of an oscillator. In exchange for a modest reduction in lifetime, we gain access to more logical quantum states in a single physical system. This could enable more efficient compilation of gates [ralph_efficient_2007, fedorov_implementation_2012] and algorithms [bocharov_factoring_2017, kiktenko_scalable_2020, gokhale_asymptotic_2019], novel techniques for quantum communication [schmidt_error-corrected_2024] and transduction [wang_passive_2024], and advantageous strategies for concatenation into an external multi-qudit code [gottesman_stabilizer_1997, gottesman_fault-tolerant_1998, campbell_magic-state_2012, campbell_enhanced_2014, Anwar_2014]. With the realization of bosonic logical qudits we also establish a platform for concatenating codes internally: by embedding a logical qubit within a bosonic logical qudit [gottesman_encoding_2001, pirandola_minimal_2008, cafaro_quantum_2012, omanakuttan_fault-tolerant_2024, gross_designing_2021, gross_hardware-efficient_2024], multiple layers of error correction could be implemented inside a single oscillator.
@@ Methods
Ethics declarations: BLB receives consulting fees from Nord Quantique. LF is a founder and shareholder of Quantum Circuits, Inc. SMG is an equity holder in, and receives consulting fees from, Quantum Circuits, Inc. The remaining authors declare no competing interests.